### v1

# Extending Audio Context for Long-Form Understanding in Large Audio-Language Models

Abstract
Large Audio-Language Models (LALMs) are often constrained by short audio context windows, even when their text backbones support long contexts, limiting long-form audio understanding. Prior work has introduced context-extension methods (e.g. YaRN) on unimodal LLMs, yet their application to LALMs remains unexplored. First, building on RoPE-based context extension, we introduce Partial YaRN, a training-free, audio-only extension method that modifies only audio token positions, leaving text positions intact to preserve the base LLM’s text capabilities. Second, we propose Virtual Longform Audio Training (VLAT), a training strategy that extends Partial YaRN into a training-time positional augmentation. VLAT simulates diverse audio lengths during training, enabling generalization to inputs far longer than those seen in training and improving robustness for long-context audio understanding. Our experiments on SALMONN and Qwen2-Audio show that Partial YaRN outperforms the original models across wide range of settings, and VLAT training strategy provides substantial improvement, achieving strong performance on long audio of unseen lengths. [^1] Code is available at: https://github.com/scb-10x/partial-yarn.
Extending Audio Context for Long-Form Understanding
in Large Audio-Language Models
Yuatyong Chaichana1,2thanks: Work done while interning at SCB 10X., Pittawat Taveekitworachai122footnotemark: 2, Warit Sirichotedumrong122footnotemark: 2, Potsawee Manakul122footnotemark: 2, Kunat Pipatanakul1thanks: Equal advising. 1SCB 10X, SCBX Group, 2Chulalongkorn University Correspondence: potsawee@scb10x.com
Figure 1: An illustration of Partial YaRN. The top plot displays low (blue) and high (red) frequency dimensions of RoPE. The bottom plot displays RoPE with Partial YaRN applied. In this example, the audio window of the low frequency dimensions are interpolated (stretched) by a factor of 1.75x to accommodate longer audio inputs, while the high frequency dimensions remain unaltered to preserve local distances and high frequency information (Peng et al., 2024). Within the Extended Audio Region, the magnitudes of every dimension is scaled to serve as a proxy for attention temperature scaling. Despite that the scaling can create a minor discontinuity in the signal, information from other dimensions can help the model disambiguate this.
1 Introduction
The success of Large Language Models (LLMs) has spurred multimodal extensions, notably large audio-language models (LALMs) [^2] Also known by other names such as Speech-aware LLMs. that pair an audio encoder with a text backbone. By aligning audio and text in a shared representation, LALMs can leverage the base LLM’s knowledge for complex audio understanding. However, practical use is constrained by short audio context windows: models such as SALMONN (Tang et al., 2024) and Qwen2-Audio (Chu et al., 2024) are typically trained on audio segments of 30s or less, and thus generalize poorly to longer inputs (Guo et al., 2025).
To address this limitation, we study the application of LLM context extension methods such as Positional Interpolation (PI) (Chen et al., 2023) and YaRN (Peng et al., 2024) for audio context extension in LALMs. To our knowledge, the application of these techniques for this specific, modality-bound use case has not been systematically explored. Applying such whole-context techniques implicitly extends the audio context window as a side effect. However, as this straightforward approach alters the positional information of the entire sequence, including text tokens; it risks degrading the sophisticated language capability of the base LLM which was pretrained solely in text.
This potential drawback motivates a more targeted strategy. We introduce “Partial YaRN”, an audio-only context extension method that modifies only the audio tokens’ positional encodings. This design enables a direct comparative study against the whole-context approaches, allowing us to investigate the tradeoff between preserving position encodings of the text modality and maintaining a globally uniform positional space.
Our experiments on both training-free and fine-tuned audio context extensions reveal that extension-based methods outperforms original models. However, the best choice between whole-context and audio-only extensions is often model-dependent without a universally superior solution. This suggests that to build truly robust long-audio-context models, we must address the core problem of generalization during the training process. Therefore, we extend Partial YaRN into a novel fine-tuning technique called “Virtual Longform Audio Training” (VLAT). Acting as a positional augmentation technique, it simulates a wide range of audio lengths during fine-tuning, teaching the model to generalize beyond the lengths present in the training dataset. Through VLAT, we obtain excellent results on longform audio of unseen lengths. Our main contributions are as follows:
  • We propose Partial YaRN, a training-free audio-context extension for LALMs that preserves the base language model’s text capabilities by leaving text positions unaltered.
  • We propose Virtual Longform Audio Training (VLAT), a training strategy that applies Partial YaRN as positional augmentation to simulate diverse audio lengths during training, improving long-context generalization.
  • We present a comparative study between existing context extension methods designed for unimodal LLMs and our proposed methods for LALMs, analyzing their performance trade-off in both training-free and fine-tuning settings.
2 Preliminary
2.1 Rotary Positional Encoding
As Transformer (Vaswani et al., 2017) utilizes attention mechanism (Bahdanau et al., 2014) which is position-invariant, explicit positional information has to be provided for the models to differentiate tokens from different positions. Our study focuses on LALMs that use Rotary Positional Encoding (RoPE) (Su et al., 2024), a widely adopted positional encoding.
RoPE formulates relative positional information through the rotation of query and key vectors. Namely, given a query $\boldsymbol{q_{m}}$ at position $m$, and key $\boldsymbol{k_{n}}$ at position $n$, we have:
$\boldsymbol{q}_{m}^{\text{T}}\boldsymbol{k}_{n}$ $=(\boldsymbol{R}_{m,\boldsymbol{\theta}}\boldsymbol{q}_{m})^{\text{T}}(\boldsymbol{R}_{n,\boldsymbol{\theta}}\boldsymbol{k}_{n})$ (1)
$=\boldsymbol{q}_{m}^{\text{T}}\boldsymbol{R}_{n-m,\boldsymbol{\theta}}\boldsymbol{k}_{n}\,$, (2)
where $\boldsymbol{\theta}$ is the set of angular frequencies (step size) of the rotation.
2.2 Context Length Extension
Transformers are usually unable to generalize well past the context window length they have seen during the training (e.g. 2048 for LLaMA (Touvron et al., 2023), 8192 for Mistral (Jiang et al., 2023)). To permit longer context length, additional fine-tuning is often required.
Position interpolation (PI).
Chen et al. (2023) proposes that instead of extrapolating RoPE further to the region unfamiliar for the trained model, interpolation is done inside the pretrained context window. Specifically, let $L$ be the original context length of the model, and $L^{\prime}$ be the target length to be extended to. Define $s=\frac{L^{\prime}}{L}$ as the extension factor. PI simply downscales the base rotational frequency by the extension factor: $\frac{\theta_{i}}{s}$. For example, if originally each positional step is rotated by $\theta_{i}=10^{\circ}$; to double the context length, the frequency (step size) can simply be halved to $5^{\circ}$. This way, the model will never have to attend to positions outside of its trained window; resulting in a more stable approach.
YaRN.
Peng et al. (2024) identifies that interpolating every dimension equally leads to the loss of local distances and high frequency information. They propose to spread the interpolation pressure by partitioning RoPE dimensions into three frequency-based groups: (1) low frequency dimensions are interpolated, (2) high frequency dimensions are solely extrapolated without interpolation to preserve high frequency information, (3) dimensions in-between get a mix of both interpolation and extrapolation. Additionally, YaRN applies temperature scaling to the logits of attention softmax:
$\operatorname{Attention}(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\operatorname{softmax}\left(\frac{\boldsymbol{Q}^{T}\boldsymbol{K}}{t\sqrt{d_{k}}}\right)\boldsymbol{V},$
where $t$ is the attention temperature.
Other methods.
Ding et al. (2024) propose LongRoPE to account for non-uniformity across RoPE dimensions through a novel evolutionary search algorithm to determine dimension-wise interpolation factors. This method is further improved in LongRoPE2 (Shang et al., 2025).
2.3 Large Audio-language Models and Their Audio Contexts
One broadly adopted paradigm of large audio-language models (LALMs) is the unified input space (Tang et al., 2024; Chu et al., 2024; Ding et al., 2025; Goel et al., 2025) which involves transforming audio inputs into the textual input space of a base LLM. By sharing the embedding space, the strong text capability and learned knowledge of the base LLM can be leveraged while simultaneously being augmented with audio understanding ability. There also exists other LALM paradigms such as cross attention (Kong et al., 2024; Ghosh et al., 2025) where the audio and text modalities are fused through cross-attention modules. This work focuses on models under the unified-input-space.
As with text context window in LLMs, LALMs also do not generalize well to audio with lengths much longer than the window seen during training (e.g. 7s for Pengi (Deshmukh et al., 2023); 30s for Qwen2-Audio (Chu et al., 2024)). Concurrent to us, Guo et al. (2025) introduces FastLongSpeech, an efficient longform LALM that uses Iterative Fusion to compress long audio representation into compact forms. Our work, in contrast, investigates an orthogonal direction of adapting the model’s positional encodings rather than the audio embeddings, and therefore is generally applicable to the existing models without requiring costly retraining.
3 Partial YaRN
To extend existing unified input LALMs to longer audio, we hypothesize that LALMs already possess a sufficient general understanding of audio and text, but the bottleneck is their unfamiliarity to audio positions beyond the range seen during audio-text training. [^3] This differs from whole-context extension in unimodal LLMs, where the total sequence length exceeds the pre-trained context window, causing an out-of-distribution problem. Here, the total sequence (audio + text) usually still fits within the original window, making the core challenge one of adapting to an unfamiliar audio length and positions, rather than extrapolating to completely unseen positions. This hypothesis suggests a training-free solution: manipulating the backbone LLM’s positional encoding. The goal is to remap positions of a long audio input into the model’s familiar audio range.
First, we study the application of YaRN (Peng et al., 2024) to LALMs, as an example of whole-context extension methods. Since the audio window is a part of the total context window, performing whole-context extension extends the audio window as a byproduct. However, this alters positional information for all tokens, including text, risking the degradation of the base LLM, which was pretrained solely in text. Motivated by the aforementioned hypothesis, we propose Partial YaRN, an audio-only extension method designed to exclusively stretch the audio window. By leaving the text’s positional encodings unaltered, this approach aims to extend [^4] The terms extend and stretch are used interchangeably. audio context while preserving the base model’s text capability.
3.1 Methodology
Inspired by PI (Chen et al., 2023) and YaRN (Peng et al., 2024) for whole-context extension in unimodal LLMs, we adapt the interpolation technique to LALMs by applying it exclusively to the audio region of the base language models.
Concretely, assuming a single audio input, let $L_{\text{audio}}$ be the original audio context length, $L_{\text{audio}}^{\prime}$ be the target length, and $p$ be the position of the first audio token. We define the positional range $p,p+L_{\text{audio}})$ as the original audio context window. It can be either a predefined location for audio input, or a dynamic region enclosed by special tokens (e.g. <speech> and </speech>). We then apply our modified YaRN technique to exclusively stretch this region to the new length $L_{\text{audio}}^{\prime}$. This creates a partially stretched positional encoding:
$\underbrace{[0,\;p)}_{\text{Text (unaltered)}}\oplus\;\underbrace{[p,\;p\!+\!L_{\text{audio}})}_{\text{{Audio} (stretched)}}\;\oplus\;\underbrace{[p\!+\!L_{\text{audio}},\;L)}_{\text{Text (unaltered)}}\,$,
where $\oplus$ denotes concatenation in the temporal axis. Note that while the final position index is denoted by $L$, the total token capacity now increases to $L+L^{\prime}_{\text{audio}}-L_{\text{audio}}$ to account for the stretched audio window of a higher token frequency.
Furthermore, we only partition the RoPE dimensions into two frequency-based groups (instead of three as in the original YaRN formulation): a low frequency group that undergoes pure interpolation, and a high frequency group that undergoes pure extrapolation. The rationale for this design is twofold:
1) Consistency: A two-group partition ensures a consistent and uniform positional encoding across the entire audio stream. YaRN’s original “in-between” group receives a mix of interpolation and extrapolation, which would cause these RoPE frequencies to not be extended to fit the entire audio. Our approach avoids this potential distortion.
2) Tuning Efficiency: This simplification reduces the tuning space of the cutoff index from two coupled variables to a single parameter.
We provide empirical validation and a more in-depth discussion of the two-group partitioning in Section 7.2 and Appendix B respectively. See Figure 1 for an example depiction of Partial YaRN.
3.2 Hyperparameters
Partial YaRN employs two hyperparameters:
1) Cutoff Dimension Index: This defines the boundary separating the RoPE dimensions into two groups. Dimensions equal or below this cutoff (low-frequency) are interpolated to stably cover longer audio context. Dimensions above it (high-frequency) are extrapolated to preserve local positional distances and high-frequency information. We default this to 0 (interpolate every dimension).
2) Attention Temperature: This controls the sharpness of attention distribution within the audio context window. A higher temperature softens the distribution, preventing attention scores from collapsing to a few tokens over long sequences. We default this to $1.0$ (no temperature scaling).
Instead of directly adding a new temperature term to the attention softmax, we follow Peng et al. (2024) and integrate the attention temperature into magnitudes of RoPE’s signals. Particularly, define $m$ and $n$ as two positions within the extended audio context window, we have the attention softmax:
$\operatorname{AS}(\boldsymbol{q}_{m},\boldsymbol{k}_{n})=\operatorname{softmax}\left(\frac{(\boldsymbol{R}_{m}\boldsymbol{q}_{m})^{\text{T}}(\boldsymbol{R}_{n}\boldsymbol{k}_{n})}{t\sqrt{d_{k}}}\right).$
Then, we invert $t$ into the two rotation matrices by scaling them with $1/\sqrt{t}$ each:
$\operatorname{AS}(\boldsymbol{q}_{m},\boldsymbol{k}_{n})\!=\!\operatorname{softmax}\!\left(\frac{\left(\frac{\boldsymbol{R}_{m}}{\sqrt{t}}\boldsymbol{q}_{m}\right)^{\text{T}}\left(\frac{\boldsymbol{R}_{n}}{\sqrt{t}}\boldsymbol{k}_{n}\right)}{\sqrt{d_{k}}}\right)\!.$
This reparameterization also handles attention between the unaltered text regions and the scaled audio regions. For instance, when a text query (unscaled) attends to an audio key (scaled by $1/\sqrt{t}$), the resulting logit naturally gets scaled by $1/\sqrt{t}$, creating a smoother temperature transition. Overall, we have temperature-scaled rotation matrices:
$\widetilde{\boldsymbol{R}}_{m}=\begin{cases}\boldsymbol{R}_{\theta,m}&\text{if }m\notin[p,p+L_{\text{audio}})\\ \frac{1}{\sqrt{t}}\boldsymbol{R}_{\theta/s,m}&\text{if }m\in[p,p+L_{\text{audio}})\end{cases}$
Notice that under the default hyperparameter values, the audio-only extension reduces from YaRN to PI, we therefore specifically name Partial YaRN with default hyperparameters: Partial PI.
4 Experimental Setup
4.1 Data
Due to the limited availability of length-tailored long audio datasets, we construct custom datasets with specific audio lengths.
YODAS2-MCQA
We source audio from the English subset of YODAS2 (Li et al., 2023b). We segment audio samples into non-overlapping segments of 1, 2, 5, and 10 minutes. For each segment, we generate five multiple-choice question-answering (MCQA) pairs using Gemini 2.0 Flash (Team et al., 2023), with four choices each. We design the generation prompt to ensure that each question focuses on different portions of the audio, and that they collectively cover the entire audio segment. The test set of each audio duration (1, 2, 5, and 10 minutes) has 750 QA pairs.
Additional details about dataset construction are provided in Appendix D.3.
Method YODAS2-MCQA
1 min 2 mins 5 mins 10 mins
[…]
Vanilla 49.01 46.13 32.93 23.47
Stretching from 30s (original context)
Whole PI 39.87 (9.14-9.14) 5.73 (40.40-40.40) 0.13 (32.80-32.80) 23.20 (0.27-0.27)
Partial PI 50.07 (+1.06+1.06) 45.60 (0.53-0.53) 12.40 (20.53-20.53) 1.20 (22.27-22.27)
Whole YaRN 54.70 (+5.69+5.69) 59.87 (+13.74+13.74) 47.20 (+14.27+14.27) 30.13 (+6.66+6.66)
Partial YaRN 57.35 (+8.34+8.34) 59.60 (+13.47+13.47) 38.53 (+5.60+5.60) 32.93 (+9.46+9.46)
Stretching from 2 mins (observed context)
Whole PI 49.01 46.13 13.07 (19.86-19.86) 23.60 (+0.13+0.13)
Partial PI 49.01 46.13 28.27 (4.66-4.66) 4.13 (19.34-19.34)
Whole YaRN 49.01 46.13 45.87 (+12.94+12.94) 38.27 (+14.80+14.80)
Partial YaRN 49.01 46.13 41.47 (+8.54+8.54) 33.45 (+9.98+9.98)
[…]
Vanilla 72.71 75.87 55.33 22.00
Stretching from 30s (original context)
Whole PI 64.50 (8.21-8.21) 7.73 (68.14-68.14) 5.87 (49.46-49.46) 8.67 (13.33-13.33)
Partial PI 72.45 (0.26-0.26) 67.73 (8.14-8.14) 40.13 (15.20-15.20) 18.40 (3.60-3.60)
Whole YaRN 73.11 (+0.40+0.40) 74.27 (1.60-1.60) 50.93 (4.40-4.40) 31.07 (+9.07+9.07)
Partial YaRN 72.72 (+0.01+0.01) 73.60 (2.27-2.27) 48.53 (6.80-6.80) 28.53 (+6.53+6.53)
Stretching from 2 mins (observed context)
Whole PI 72.71 75.87 46.93 (8.40-8.40) 21.20 (0.80-0.80)
Partial PI 72.71 75.87 59.60 (+4.27+4.27) 45.73 (+23.73+23.73)
Whole YaRN 72.71 75.87 60.27 (+4.94+4.94) 47.60 (+25.60+25.60)
Partial YaRN 72.71 75.87 60.40 (+5.07+5.07) 48.00 (+26.00+26.00)
Table 1: Training-free audio context extension performance (Accuracy). We compare extension methods under two strategies: stretching from the original 30s context and the empirically observed 2mins context. We see that stretching from the observed 2mins context proves significantly more effective. Comparing between the two top approaches: Whole and Partial YaRN, we find their performance to be closed, suggesting that preserving the positional information of text tokens might not be of utmost importance, at least in this MCQA setting of ours.
4.2 Models
We select two widely-used open-weights LALMs: SALMONN (Tang et al., 2024) and Qwen2-Audio (Chu et al., 2024). Both models utilize Whisper-encoder (Radford et al., 2023) as audio encoder, which is limited to processing audio of up to 30s. We handle longer audio inputs by segmenting them into non-overlapping 30s chunks, where each chunk is encoded independently.
The number of audio representations differs considerably between these models. For each 30s chunk on 50Hz Whisper-encoder, SALMONN generates a sequence of 88 audio tokens (with Q-Former), whereas Qwen2-Audio produces a sequence of 750 tokens (with a 2x downsampling).
4.3 Baseline Methods
We evaluate Partial PI (default hyperparameters) and Partial YaRN (tuned hyperparameters) against the following methods:
1) Vanilla: Unmodified base models. Long audio inputs are passed directly without any manipulation to their RoPE.
2) Whole Position Interpolation (PI) (Chen et al., 2023): Every RoPE dimension is uniformly interpolated across the whole context window. This serves as the primitive whole-context baseline.
3) Whole YaRN (Peng et al., 2024): Frequency-based RoPE interpolation with attention scaling, applied to the whole context window. This serves as the primary whole-context baseline.
These three baselines require no tuning as Vanilla and Whole PI don’t utilize any hyperparameter, and YaRN has predefined cutoffs and a closed-form formula for obtaining temperature.
Reference models.
We also report performance of two proprietary models capable of long audio context: GPT-4o (OpenAI, 2024) and Gemini 2.0 Flash (Team et al., 2023), in order to provide a broader perspective of LALMs and to validate the quality of our dataset [^5] The model versions are gpt-4o-mini-2024-07-18 and gemini-2.0-flash-001 respectively. Note that Gemini was also used for generating our YODAS2-MCQA dataset..
Vanilla Inference Partial PI Inference
Training Method 2 mins 5 mins 10 mins 2 mins 5 mins 10 mins
Vanilla Fine-tuning 97.60 89.07 32.76 96.40 89.78 75.56
Virtual Longform 97.60 (+0.00+0.00) 91.42 (+2.35+2.35) 75.11 (+42.35+42.35) 96.87 (+0.47+0.47) 91.91 (+2.13+2.13) 81.73 (+6.17+6.17)
Table 2: Virtual Longform Audio Training (VLAT) enables strong generalization to unseen audio lengths (Accuracy). We compare standard Vanilla fine-tuning and our proposed VLAT on Qwen2-Audio, evaluated with (Partial PI Inference column group) and without (Vanilla Inference column group) Partial PI during inference.
5 Main Results
In this section, we comparatively study the effect of whole-context and audio-only context extensions on both the training-free and fine-tuning settings for audio segments up to 10mins.
5.1 Training-free Audio Context Extension
First, we extend the audio context length without training. We report the results in Table 4.1, showing that: no single method—whole-context or audio-only—is universally superior. Instead, the performance varies between the different models and the degree of extension. More importantly, we observe that both models generally retain consistent performance up to 2mins, suggesting that their innate audio context lengths are longer than just 30s, possibly a result from multi-audio training.
Extension of SALMONN.
When stretching from the 30s window, extension methods provide substantial performance gains compared to the Vanilla baseline. We observe that Partial YaRN shows the strongest performance on the 1min and 10mins settings, achieving 57.35% and 32.93% accuracies. On 2mins, both Whole and Partial YaRN perform similarly. However, Whole YaRN outperforms Partial YaRN at 5mins by a large margin. Possible reasons for this performance flip are (1) higher expressivity and less compression pressure from Whole YaRN’s additional “in-between” dimension group, outweighs the preservation of the base language capability in this 5mins setting, or (2) noise in the hyperparameter tuning process of Partial YaRN, which is an inherent drawback.
Extension of Qwen2-Audio.
In contrast to SALMONN, Qwen2-Audio demonstrates different characteristics. The Vanilla baseline is more robust, outperforming all stretch-from-30s methods at 2mins. Both Whole and Partial YaRN struggle to provide significant gains and even underperform the baseline at moderate lengths (2 and 5mins) when stretching from the original 30s context; however, they manage to attain considerable gains on the 10mins setting. This indicates that Qwen2-Audio likely possesses a strong audio-length generalization capability or longer innate audio context window. For this model, simply applying an extension method from the base context does not guarantee an improvement. This suggests that the model’s high intrinsic audio-length generalizability is superior to the extension methods that inevitably cram the positional information as a side effect.
Extending from observed audio context.
As previously mentioned in the beginning of this section that we can observe both model’s innate audio contexts to be around 2mins instead of just 30s, here we experiment with extending from this observed audio context of 2mins to the target lengths. For both SALMONN and Qwen2-Audio, performance on the long audio improves substantially when the interpolation is anchored from a 2mins context instead of the original 30s. Notably on Qwen2-Audio, this strategy boosts Partial YaRN’s performance at 10mins from 28.53% to a 48.00%, an absolute improvement of almost 20%, yielding a final performance 26% higher than the Vanilla baseline. This finding suggests that determining the extension by observing the vanilla performance may be a more critical factor than the choice between whole-context or audio-only methods, especially on large extension ratios.
5.2 Fine-tuning for Audio Context Extension
Next, we study training-based extension methods. Keeping other components frozen, we fine-tune the base language model of Qwen2-Audio using a single epoch LoRA (Hu et al., 2022) of rank 8 on the YODAS2-MCQA training set of the corresponding length. Two LoRA settings are explored: adapting only the query projection (q), and adapting query, key, value, and output projections (qkvo). Due to the high cost of tuning Partial YaRN’s hyperparameters during training, we only use the default configuration (Partial PI) for the audio-only method. Whole PI is omitted due to its poor performance as previously observed.
The results are presented in Table 3, showing that fine-tuning with a context extension method dramatically outperforms the vanilla fine-tuning baseline, especially at longer audio lengths. In the qkvo setting, both Whole YaRN (83.47%) and Partial PI (83.07%) achieve an absolute performance gain of around 19% over the Vanilla baseline (64.93%) at 10mins. This highlights the benefit of integrating a context extension method into the fine-tuning process for more better context extension.
Comparing the whole-context and audio-only approaches, we find that their performance is generally competitive. While Whole YaRN triumps when only adapting the query projection weights, suggesting it’s ease of adaptation. Their performance converges under standard LoRA practice which adapts qkvo. This result confirms that both whole-context and audio-only methods are highly effective strategies for fine-tuning.
Adapt Method 2 mins 5 mins 10 mins
q Vanilla 95.07 78.93 21.60
Whole YaRN 96.53 88.53 73.73
Partial PI 96.13 83.73 69.20
qkvo Vanilla 96.13 85.20 64.93
Whole YaRN 97.47 93.60 83.47
Partial PI 98.00 93.33 83.07
Table 3: Finetuned audio context extension performance (Accuracy). We finetune Qwen2-Audio on YODAS2-MCQA, under two LoRA settings: adapting only q, and adapting qkvo. Each model is finetuned and evaluated using the same extension method.
6 Virtual Longform Audio Training: Generalizing Beyond the Training Data
Our previous results show that while the training-free methods are beneficial, their effectiveness is often model and audio-length dependent. Based on this, we propose to fine-tune LALMs with Partial YaRN repurposed as a positional augmentation technique we call "Virtual Longform Audio Training" (VLAT).
6.1 Methodology
The goal of VLAT is to simulate and expose models to audio context windows of diverse lengths during training. Let $L_{\text{data}}$ be the actual length of a training audio sample. For each sample, we obtain a “virtual” source length, $L_{\text{virt}}$, as the model’s default audio context length times a factor randomly sampled from the range [1, 5, 10, 15, 20, 25]x. We then apply Partial YaRN [^6] Whole YaRN is not used here, as we observe that it diverges quickly when training under the VLAT framework. to stretch or compress the $L_{\text{virt}}$-long positional window to $L_{\text{data}}$.
For instance, consider an audio sample with $L_{\text{data}}=2$mins, and we draw a $L_{\text{virt}}$ of 10mins, Partial YaRN will compress a 10mins context window to 2mins for this audio. This process effectively simulates the 2mins audio to be 10mins long virtually, thereby familiarizing the model to longer audio context and improving its ability to generalize to genuinely long audio at inference time.
Randomizing positional encoding has previously been explored in Ruoss et al. (2023), where positional indices are randomly downsampled to let the model see larger positional values. Rather than using $\{1,2,\dots,N\}$ for an length-$N$ input, they use a randomly subsampled values such as $\{1,4,\dots,L\}$ with $L>N$ instead. In contrast, VLAT differs in employing interpolation inside the context window to create a dense, continuous space of positions, unlike sparse integer subsampling. It is also bidirectional, teaching the model through both compressed and stretched contexts. Furthermore, VLAT is a targeted, modality-bound method that modifies only the audio tokens, preserving the base LLM’s original textual space.
6.2 Evaluation
We fine-tune Qwen2-Audio on a 2mins YODAS2-MCQA training set [^7] To prevent data leakage, the train/test splits were performed at the video level, ensuring no source material overlaps between the training set and any of the test sets. using two methods: a standard Vanilla fine-tuning and VLAT. We employ 10-epoch qkvo LoRA for the adaptations. We then evaluate the models from both training methods under two different extensions at inference: Vanilla and Partial PI. See Appendix E.1 for additional detail on VLAT configuration.
6.3 Results
As shown in Table 2, VLAT greatly improves generalization to unseen audio lengths. When evaluated with vanilla inference, the model trained with Virtual Longform shows a dramatic improvement on 10mins audio, increasing accuracy from 32.76% to 75.11%, closing in on the previous 1-epoch direct-fine-tuning result. This highlights the effectiveness of the training technique for training LALMs to generalize far beyond the audio lengths present in their training data, mitigating a key bottleneck in the development of robust long-audio models.
Furthermore, as can be seen in "Partial PI Inference" columns, VLAT’s benefit is complementary to inference-time extension. When Partial PI is applied during evaluation, the VLAT-trained model again outperforms its vanilla-trained counterpart, achieving the highest overall 10mins performance of 81.73%. This indicates that VLAT and inference-time extension are compatible strategies, and their combination yields the most robust long-context performance.
7 Analysis
7.1 Ablation Study
Here, we conduct an ablation study to isolate and analyze the individual contributions of the two key components in Whole YaRN (Peng et al., 2024) and Partial YaRN: (1) frequency grouping and (2) attention temperature scaling. We note that by removing both of these components, the methods converge to Whole PI (Chen et al., 2023) and Partial PI. The results are presented in Table 7.1.
We can see that removing either the frequency grouping or the attention temperature generally leads to a large drop in performance, especially on higher extension ratios. Both the frequency partitioning and attention temperature scaling are crucial for robustly extending the audio context length of LALMs. Their combined application within the Whole and Partial YaRN frameworks generally yield the most effective and stable performance.
Partial Whole
Method 2 mins 10 mins 2 mins 10 mins
[…]
YaRN 59.60 32.93 59.87 30.13
- Freq grouping 56.93 4.53 23.33 29.20
- Attn temp 48.13 23.87 41.07 23.20
PI 45.60 1.20 5.37 23.20
[…]
YaRN 73.60 28.53 73.47 25.33
- Freq grouping 66.00 28.13 17.07 9.87
- Attn temp 73.47 18.80 75.07 22.80
PI 67.73 18.40 7.73 8.67
Table 4: Ablation of frequency grouping and attention temperature components (Accuracy). Removing either the frequency grouping or the attention temperature leads to a large drop in performance.
7.2 Partial YaRN with Three-group Frequency Partitioning
For the Partial YaRN method, we modify the original 3-group frequency partition from original YaRN to a 2-group method as discussed in Section 3.1 and Appendix B. Here we conduct a direct empirical comparison between the 2-group and 3-group approaches.
For the 3-group baseline, we implement Partial YaRN using YaRN’s original partitioning scheme and its recommended hyperparameters. We compare this against our proposed 2-group Partial YaRN using the hyperparameters from our main experiments (Section 5.1). We present the results in Table 5.
YODAS2-MCQA
Model Freq Grouping 2 mins 5 mins 10 mins
SALMONN 2-Group 59.60 38.53 32.93
3-Group 56.00 41.87 25.46
Qwen2-Audio 2-Group 73.60 48.53 28.53
3-Group 72.00 20.00 16.53
Table 5: Performance Comparison of Two-Group vs. Three-Group Partitioning (Accuracy). We compare our Partial YaRN with 2-group frequency partitioning used throughout the work, against a 3-group version.
We see that our 2-group partitioning broadly outperform 3-group in our modality-specific scenario. This is most evident with the Qwen2-Audio model, where the 3-group method suffers a large performance drop on 5mins and 10mins audio, with an accuracy gap of 28.53% and 12.00% respectively. Overall, this result validates the effectiveness and reliability of our proposed 2-group scheme used in Partial YaRN.
8 Conclusion
This work addressed the challenge of extending audio context windows of LALMs by studying the application of LLM’s whole-context extension methods, and proposing a training-free audio-only extension Partial YaRN and the VLAT training strategy. We demonstrate the effectiveness of extension-based methods over non-extension baselines across various long-audio settings. Leveraging our finding that SALMONN and Qwen2-Audio retain consistent performance up to 2mins audio window, we achieved an even stronger performance by extending from the 2mins window instead of 30s. Later, we showed that fine-tuning through VLAT helps the models to generalize to audio contexts far exceeding the length of the training data, offering a robust and data-efficient pathway to develop LALMs with better longform understanding. Future work could explore a similar modality-bound extension strategy for video models.
Limitations
Evaluation scope.
Our experiments are conducted exclusively on the multiple-choice question answering (MCQA) tasks. We made this choice to ensure a precise and unambiguous measurement of the model performance under different audio lengths. This mitigates the confounding variables inherent in evaluating open-ended generation, such as judging the semantic equivalence or stylistic alignment. However, a key tradeoff of this focused evaluation is its limited ability to assess the nuanced generative and language modeling capabilities of the base LLM. Consequently, while our results validate the model’s ability to retrieve information from audio, they do not fully test the importance of text preservation of the audio-only methods like Partial YaRN. Furthermore, our study does not assess performance across other long-audio task formats found in benchmarks such as (Guo et al., 2025; Ahia et al., 2025; Goel et al., 2025).
Hyperparameter tuning burden.
Unlike the original YaRN which has generally good predefined cutoff dimensions and a closed-form formula for determining the attention temperature, our proposed Partial YaRN requires tuning of these two hyperparameters, thereby introducing additional computational burden.
Acknowledgments
YC would like to thank Ekapol Chuangsuwanich for insightful discussion on the earlier version of this work. He also thanks Thanapat Trachu for valuable guidance on the illustration of Partial YaRN.
Appendix A Extended Related Works
This section discusses additional related works that aren’t addressed in Section 2.
Positional encodings.
Conceived along with the introduction of transformers (Vaswani et al., 2017), the original absolute positional encoding utilizes sinusoidal waves of varying frequencies to uniquely represent absolute position of the tokens, and is incorporated into the model by adding directly into the input sequence. Later works (Devlin et al., 2019; Lan et al., 2020; Clark et al., 2020) propose to use trainable vectors for representing absolute positional information, where token at each position is summed with a learned vector, allowing the models to learn the optimal positional encoding for themselves. However, it cannot extrapolate to sequences longer than the maximum length it was trained on. Shaw et al. (2018) introduces to encode the relative position of the token instead, where the relative distance between the query and key is captured and injected into every attention layer. Dai et al. (2019) modifies the decomposed form of the attention equation, and introduces additional trainable location-agnostic key vectors to distinguish between content-based and location-based querying behaviors. Raffel et al. (2020) groups the relative distances into buckets, and associate each bucket with a learnable scalar for adding into the attention logits. Simplifying the encoding scheme and reducing trainable parameters. TUPE (Ke et al., 2021) uses separate linear projections for the word and positional information. ALiBi (Press et al., 2022) simply adds linear biases with a learnable slope into the attention scores, the bias grows linearly with the relative distance between the tokens. This method generally imposes a strong locality preference by encouraging models to attend more to the nearby tokens, while discouraging them from attending to tokens that are farther away. Several later works extend RoPE for multimodal modeling (e.g. Multimodal RoPE Bai et al. (2025), VideoRoPE (Wei et al., 2025), VRoPE (Liu et al., 2025)).
Large video-language models.
Video input is similar to the audio modality in that they’re both naturally continuous and can be of arbitrary length. Many Large Video-Language Models (LVLMs), such as Video-LLaMA (Zhang et al., 2023b) and Video-ChatGPT (Maaz et al., 2024), adopt architectural paradigm similar to the unified input space LALMs we study, where video frames are embedded and projected into a base LLM’s input space. However, the context length problem is even more exacerbated in the video domain due to the high dimensionality and large number of tokens needed to represent even short videos. To improve efficiency, the LVLM community has largely focused on input compression techniques like query aggregation (Zhang et al., 2023b; Li et al., 2023a; Song et al., 2024), frame pooling (Luo et al., 2023; Maaz et al., 2024), and feature merging (Weng et al., 2024). These strategies primarily aim to reduce the number of tokens before they enter the LLM. Our work offers a complementary perspective: instead of altering the input, we manipulate the positional space within the model to accommodate a longer sequence. The principles of Partial YaRN and Virtual Longform Audio Training could therefore be potentially applicable to the video domain as well, hopefully providing an orthogonal path to achieving long-context understanding in synergy with existing compression methods.
Appendix B Elaboration on Two-group Partitioning of Partial YaRN
The decision to simplify the original three-group frequency partition of YaRN (Peng et al., 2024) into a two-group system for Partial YaRN is a deliberate design choice aimed at ensuring positional consistency across the entire extended audio stream. The original YaRN formulation partitions RoPE dimensions into: 1) a low-frequency group that is purely interpolated, 2) a high-frequency group that is purely extrapolated, and 3) an "in-between" group that receives a mix of both, with a ramped interpolation factor.
While effective for whole-context extension, this "in-between" group becomes problematic in our partial, audio-only extension scenario. Consider a 2mins audio input, which requires a 4x extension of the model’s 30s original audio context.
  • The high-frequency dimensions, when extrapolated (a 1x factor), correctly preserve local relative distances, which is desirable.
  • The low-frequency dimensions, responsible for quasi-absolute positioning, are fully interpolated by a 4x factor. This is crucial as it maps the entire 2mins audio into the model’s familiar positional range (original audio context).
  • The "in-between" dimensions, however, would receive interpolation factors between the 1x and 4x (e.g. 2x). This means that for these specific dimensions, the model’s original context window is stretched to cover only the first minute of the 2mins audio. The second minute would lie in an extrapolated, out-of-distribution region.
This dimensional inconsistency risks creating a biased or distorted representation of the latter half of the audio. The model would receive conflicting signals about whether a token is "inside" or "outside" its familiar context.
By simplifying to a two-group partition, we circumvent this issue entirely. Every dimension is forced into a binary choice: either it is fully interpolated to cover the entire audio duration (low-frequency), or it is fully extrapolated to preserve local structure (high-frequency). This ensures that the model maintains a consistent positional understanding across all dimensions for the entirety of the audio input, avoiding the representational bias that partial coverage would introduce.
See Appendix 7.2 for direct performance comparison between 2 and 3-group frequency partitioning versions of Partial YaRN.
Appendix C Implementation of Partial YaRN via position_ids Manipulation
A key implementation detail of our work is how we apply the positional stretching to only a subsection of the input sequence. While the original YaRN implementation modifies the RoPE frequency variables ($\boldsymbol{\theta}$), this approach isn’t straightforward for a partial modification. Instead, our implementation of Partial YaRN directly manipulates the position_ids and multiply it with the unmodified frequencies. This section details the implementation our method.
Modifying positional indices.
In most Transformer architectures, including the models used in this work, the RoPE mechanism computes embeddings based on an input tensor called position_ids. This tensor is simply a sequence of integers representing the absolute position of each token (i.e. [0, 1, 2, ..., N-1]). Our method leverages this by modifying a section this tensor. The process is as follows:
  1. 1.
    Identify the Audio Region: First, we identify the start and end indices of the audio tokens within the position_ids tensor. This creates three distinct segments: the leading text positions (unaltered), the audio positions (to be stretched), and the trailing text positions (unaltered relative to each other).
  2. 2.
    Stretch the Audio Positions: We replace the original sequence of integer positions for the audio segment with a new sequence generated via torch.linspace. For example, if the original audio positions were [p, p+1, ..., p+base_L_audio-1] and the new audio input has L_audio tokens, we generate L_audio new positions by interpolating within the original range: torch.linspace(start=p, end=p+L_audio-1, steps=L_audio). This effectively “stretches” the familiar positional space to accommodate the longer audio input.
  3. 3.
    Concatenate: The final position_ids tensor is constructed by concatenating the leading text positions, the new stretched audio positions, and the trailing text positions. This new position_ids is now ready to be use for subsequent construction of partially stretch RoPE signal for the relevant dimensions.
Computational overhead.
This implementation introduces negligible computational overhead in practice. The only additional computations are incurred during the construction of the position_ids tensor, which involves a few tensor slicing operations, a single torch.linspace call, and a final concatenation. These are highly optimized, vectorized operations that constitute a minuscule fraction of the total computation in a full forward pass. Crucially, this precomputation is a one-time cost per input sample, applied only during the initial processing of the prompt that contains the long audio. During the subsequent autoregressive generation steps, where tokens are generated one at a time, our method adds no overhead whatsoever. The model reverts to the standard process of incremental positional stepping, just as it would without any context extension. Therefore, the cost of enabling partial context extension is a small, fixed precomputation that does not impact the per-token generation latency.
Appendix D Additional Experimental Details
D.1 Training and Inference
Environment.
All fine-tunings and inferences are conducted using PyTorch (Paszke et al., 2019) and HuggingFace’s Transformers (Wolf et al., 2020) libraries on a single Nvidia H100 GPU. When available, we utilize Flash-Attention 2 (Dao, 2024) to speed up the processes and reduce VRAM consumptions. For experiments on very long sequence, we use gradient checkpointing to circumvent out-of-memory issue.
fine-tuning hyperparameters.
We use LoRA (Hu et al., 2022) of rank 8 for fine-tuning through HuggingFace’s PEFT library (Mangrulkar et al., 2022). With alpha set to 16, and a small LoRA dropout of 0.5%. Models are trained with AdamW optimizer (Loshchilov and Hutter, 2019), learning rate is set to 5e-5, a batch size of 8 samples, and gradient norm clipped to 1.0 (Pascanu et al., 2013).
D.2 Partial YaRN’s Hyperparameter Tuning
Recall that Partial YaRN has two hyperparameters, cutoff dimension and attention temperature. For the training-free experiment (Section 5.1), we jointly sweep the cutoff dimension from $\{56,48,40,32,24,16,8\}$, and the temperature from between 0.5 and 1.6 with a step size of 0.1. See Appendix E.2 for tuning results and sensitivity.
D.3 Dataset Construction
We specifically use a portion the en129888https://huggingface.co/datasets/espnet/yodas2/tree/main/data/en129 subset of the YODAS2 dataset (Li et al., 2023b) (CC-BY-3.0) through HuggingFace’s Datasets library (Lhoest et al., 2021). The dataset is then split into train, validation, and test sets. We guarantee no test leakage in a way that if a dialogue is in the test set, then all the other dialogues from the same video are enforced to be in the test set as well.
YODAS2 MCQA inference prompt Your answer MUST be in the format of "(NUMBER) STATEMENT". For example, if the answer was (4) A pen, you would ONLY output "(4) A pen". Do NOT include any other text. Using information from all the given audios. {question}
Figure 2: Prompt used for inferencing on YODAS2-MCQA.
To obtain specific audio length, we chunk the long audio in the dataset to 1, 2, 5, and 10 minutes. Last, incomplete chunk of each long-audio is also included as well. Each of the chunk will be used to generate 5 MCQA pairs via Gemini 2.0 Flash (Team et al., 2023). We provide the full MCQA generation prompt in Figure 7.
Appendix E Further Experimental Results
E.1 Different VLAT’s Virtual Length Sampling Strategies
The effectiveness of Virtual Longform Audio Training (VLAT) depends on the strategy used to sample the virtual length during training. For generality, here we define the virtual length as a produce of true audio context length (i.e. 30s) and the virtual factor. In our main experiment Section 6, we used a strategy of randomly sampling a factor from the range [1x, 5x, 10x, 15x, 20x, 25x]. To understand the sensitivity of our results to this choice, we compare our default strategy against four alternatives, training a Qwen2-Audio model on each strategy and evaluating on the 10mins audio task.
The virtual factor sampling strategies are:
  1. 1.
    Default: Uniform random sampling from $\{1.0,5.0,10.0,15.0,20.0,25.0\}$. This cover the audio context length of 30s (1.0x) to 12.5mins (25.0x).
  2. 2.
    Dense Sampling: Random sampling from a discrete set of 100 uniformly spaced points within the default strategy’s range.
  3. 3.
    Very Dense Sampling: Random sampling from a discrete set of 1000 uniformly spaced points within the default strategy’s range.
  4. 4.
    Limited Range: Uniform random sampling from $\{1.0,5.0,10.0\}$. This factor set covers audio context length of at most 5mins (10.0x).
  5. 5.
    Fixed High-Factor: No randomization; the scale factor is always fixed at 20.0x.
We report the result in Table 6.
YODAS2-MCQA
Virtual Length Strategy 2 mins 5 mins 10 mins
Default 96.87 91.91 81.73
100 Sampling Points 95.91 89.25 79.78
1000 Sampling Points 96.18 87.69 79.60
Limited Range 98.00 93.55 89.20
Fixed to 20.0x 89.60 68.71 50.31
Table 6: VLAT performance with different sampling strategies. All models are trained on 2mins audio and evaluated with Partial PI inference. The results are averaged across three random seeds.
Analysis.
The experiment offers several insights into different VLAT sampling strategies: (1) Fixed High-Factor (always 20.0x) strategy performs the worst by a large margin, confirming that exposure to a variety of positional ranges is crucial for robust generalization, rather than training for just a single fixed target length. (2) Default strategy which uses a coarse set of 6 sampling points outperforms both the Dense (100 points) and Very Dense (1000 points) strategies. This suggests that finegrained sampling of virtual factors provides no additional benefit, and that a relatively small but diverse set of factors is sufficient for training length-robust models.
The most surprising result comes from the Limited Range strategy, which achieves the highest performance across all evaluation lengths, despite its virtual lengths not extending beyond the 5mins mark (10.0x factor). While this demonstrates a powerful capacity for extrapolation, the mechanism behind this superior performance is not immediately clear. It may suggest that randomization on extremely long virtual lengths could introduce instability to the training, which could in turn explain why Whole YaRN diverged quickly under VLAT framework; or that the specific distribution of factors in this limited set is coincidentally optimal for our test data. We left the more in-depth study of this behavior as future work.
Given this uncertainty, we retained the Default strategy for our main experiments in Section 6. Its factor range intuitively and completely covers the lengths seen during evaluation, making it a more principled and safer choice.
E.2 Hyperparameter Sensitivity of Partial YaRN
Here we analyze the effect and sensitivity of different Partial YaRN’s hyperparameter configurations. We report validation accuracies w.r.t. different cutoff dimension indices and attention temperatures, for the Qwen2-Audio and SALMONN models.
E.2.1 Sensitivity on Qwen2-Audio
Figure 3 displays the hyperparameter sensitivity of Partial YaRN on Qwen2-Audio when extending from the original 30s audio context window. Figure 4 shows the sensitivity when extending from the observed audio context window of 2mins.
Sensitivity trend.
We observe the trend of optimal hyperparameter configurations to move from the bottom left (high cutoff dimension index, temperature close to 1.0) toward the top right of the heatmaps (lower cutoff dimension index, higher temperature) as the audio length increases.
Hyperparameter recommendation.
This trend informs the following recommendation when using Partial YaRN. For moderate context extensions (e.g., 2x to 4x), it is advisable to start with a higher cutoff dimension index and a temperature value close to 1.0. For longer extensions, a lower cutoff dimension index with a considerably higher temperature is likely to yield better results.
Plausible explanation of this trend.
First, the attention temperature controls the distribution of attention scores. In very long sequences with thousands of tokens, the standard softmax function can lead to an uneven attention score distribution, such as attending to only a few tokens while ignoring the vast majority of the context. Raising the temperature flattens this distribution, encouraging the model to maintain a broader attention pattern across the entire long audio stream and preventing "attention collapse".
Second, the cutoff dimension index directly controls the "compression pressure" of the position encoding extension. A high cutoff index is expressive enough for shorter extensions while retaining majority of the original, detailed positional information. On the other hand, a lower cutoff index (higher compression) is required to map a much larger range of positions into the smaller familiar audio region.
E.2.2 Sensitivity on SALMONN
Figure 5 displays the hyperparameter sensitivity of Partial YaRN on SALMONN when extending from the original 30s audio context window.
Sensitivity trend.
For this model, we observe that the optimal hyperparameter sets are generally contained in the low temperature region (0.5 - 0.7), with the cutoff dimension index moving from low to higher ones as the audio length grows.
Hyperparameter recommendation.
From the observed trend, it’s advisable to set the attention temperature to $0.6(\pm 0.1)$, adopt a low cutoff index for shorter audio length, and higher index for longer audio inputs.
Plausible explanation of this trend.
The vast distinction between SALMONN and the previously discussed Qwen2-Audio lies in the their audio representation lengths—SALMONN only uses 88 tokens for every 30s audio vs. Qwen2-Audio’s 750 tokens. This, coupled with the model’s preference of low attention temperature, suggests that letting the model focuses on a relevant local region outweighs the need to spreading the attention distribution as in Qwen2-Audio.
Figure 3: Hyperparameter Sensitivity of Partial YaRN on Qwen2-Audio (extending the original 30s audio context).
Figure 4: Hyperparameter Sensitivity of Partial YaRN on Qwen2-Audio (extending the observed 2mins audio context).
Figure 5: Hyperparameter Sensitivity of Partial YaRN on SALMONN (extending the original 30s audio context).
Figure 6: Hyperparameter Sensitivity of Partial YaRN on SALMONN (extending the observed 2mins audio context).
YODAS2 MCQA generation prompt You are an expert in creating high-quality Multiple-Choice Question Answering (MCQA) datasets for speech understanding. Your task is to generate QA pairs based on audio segments. **Input:** You will be given a segment of an audio recording. **Task:** For each given transcript, generate **5 distinct 4-Choice Question-Answer (QA) pairs** that each has 1 correct and 3 incorrect choices (so four choices in total), and are factually grounded in the content of the transcript. The questions should cover different aspects of the audio segment, such as: * **Who:** Identify speakers or people mentioned. * **What:** Describe actions, objects, or events. * **When:** Determine times or dates mentioned. * **Where:** Locate places or locations. * **Why/How:** Explain reasons or processes. * **Details:** Ask for specific details mentioned in the audio segment. * **Purpose/Intention:** Infer the speaker’s intent. * **Summary:** Condense information from the audio segment. * **Count:** How many people, men or women in the audio segment. * **Opinion/Sentiment:** What opinions or sentiment is expressed. * **Negation:** Understand what is NOT being said or done. **Output:** Your output should be a JSON formatted string. The JSON should have the following structure: "response": [ "question": "Question 1", "correct": "Correct answer", "incorrect1": "Incorrect answer1", "incorrect2": "Incorrect answer2", "incorrect3": "Incorrect answer3", "question": "Question 2", "correct": "Correct answer", "incorrect1": "Incorrect answer1", "incorrect2": "Incorrect answer2", "incorrect3": "Incorrect answer3", "question": "Question 3", "correct": "Correct answer", "incorrect1": "Incorrect answer1", "incorrect2": "Incorrect answer2", "incorrect3": "Incorrect answer3", "question": "Question 4", "correct": "Correct answer", "incorrect1": "Incorrect answer1", "incorrect2": "Incorrect answer2", "incorrect3": "Incorrect answer3", "question": "Question 5", "correct": "Correct answer", "incorrect1": "Incorrect answer1", "incorrect2": "Incorrect answer2", "incorrect3": "Incorrect answer3" ] Important Guidelines: Accuracy: Ensure the answers are directly supported by content of the audio. Do not introduce information that is not present in the audio. Relevance: The questions should be relevant to the content of the audio segment. Diversity: The 5 QA pairs should be diverse in the type of question asked (who, what, when, where, why, how, detail, etc.). Try to use each question type at least once, if applicable to the transcript. Clarity: The questions and answers should be clear, concise, and grammatically correct. Completeness: The answers should fully address the question and provide sufficient information. Conciseness: The answers should be as concise as possible while still being accurate and complete. Coverage: The questions should collectively cover the entire audio segment. Aim for diverse coverage for each question. Prioritize the most salient and important information. Avoid Ambiguity: Ensure the questions are not ambiguous and have a clear, single answer. Plausibility of Incorrect Answers: The incorrect answers should be plausible but wrong, and related to the context of the transcript. Don’t make them obviously incorrect or nonsensical. Use information or wording from the transcript to make them believable distractors. JSON Validity: The output must be a valid JSON string. Double-check the formatting and escaping of special characters. Context-Awareness: While each audio segment should be treated independently, the types of questions should logically relate to the context of human conversations. Try to create questions which someone would naturally ask. Answer Length Consistency: Try to make the length of the correct and incorrect answers as similar as possible. This makes it more difficult to guess the answer based on length alone. Example: Input Transcript: "Okay, so the meeting is scheduled for next Tuesday at 2 PM in Conference Room A. We’$l$ be discussing the Q3 marketing report and the proposed budget changes." Output: "response": [ "question": "When is the meeting scheduled?", "correct": "Next Tuesday at 2 PM", "incorrect1": "This Friday at 10 AM", "incorrect2": "Next Wednesday at 3 PM", "incorrect3": "Today at 4 PM" , "question": "Where will the meeting take place?", "correct": "Conference Room A", "incorrect1": "Conference Room B", "incorrect2": "The Main Auditorium", "incorrect3": "The Executive Boardroom" , "question": "What marketing report will be discussed?", "correct": "The Q3 marketing report", "incorrect1": "The Q2 marketing report", "incorrect2": "The annual marketing report", "incorrect3": "The preliminary marketing report" , "question": "Besides the marketing report, what else will be discussed?", "correct": "The proposed budget changes", "incorrect1": "The sales figures", "incorrect2": "The employee satisfaction survey", "incorrect3": "The new product launch" , "question": "What is the purpose of the meeting?", "correct": "To discuss the Q3 marketing report and budget changes", "incorrect1": "To plan the annual company picnic", "incorrect2": "To interview new job candidates", "incorrect3": "To celebrate the company’s anniversary" ] Now, generate 5 QA pairs in the specified JSON format for the input audio:
Figure 7: Prompt for generating multiple-choice question answering pairs from YODAS2.

### v2

@@
 YaRN) on unimodal LLMs, yet their application to LALMs remains unexplored.
-First, building on RoPE-based context extension, we introduce Partial YaRN, a training-free, audio-only extension method that modifies only audio token positions, leaving text positions intact to preserve the base LLM’s text capabilities.
+First, building on RoPE-based context extension, we introduce Partial YaRN, a training-free, modality-decoupled extension method that modifies only audio token positions, leaving text positions intact to preserve the base LLM’s text capabilities.
 Second, we propose Virtual Longform Audio Training (VLAT), a training strategy that extends Partial YaRN into a training-time positional augmentation.
-VLAT simulates diverse audio lengths during training, enabling generalization to inputs far longer than those seen in training and improving robustness for long-context audio understanding.
-Our experiments on SALMONN and Qwen2-Audio show that Partial YaRN outperforms the original models across wide range of settings, and VLAT training strategy provides substantial improvement, achieving strong performance on long audio of unseen lengths.
-[^1] Code is available at: https://github.com/scb-10x/partial-yarn.
+VLAT simulates diverse audio lengths during training, enabling generalization to inputs far longer than those seen in training.
+Our experiments on SALMONN and Qwen2-Audio confirm that Partial YaRN outperforms the original models across wide range of settings, and VLAT provides substantial performance improvement on long audio of unseen lengths.
+[^1] Code and dataset are available at: https://github.com/yophis/partial-yarn.
 Extending Audio Context for Long-Form Understanding
@@
 Yuatyong Chaichana1,2thanks: Work done while interning at SCB 10X., Pittawat Taveekitworachai122footnotemark: 2, Warit Sirichotedumrong122footnotemark: 2, Potsawee Manakul122footnotemark: 2, Kunat Pipatanakul1thanks: Equal advising.
-1SCB 10X, SCBX Group, 2Chulalongkorn University Correspondence: potsawee@scb10x.com
+1SCB 10X, SCBX Group, 2Chulalongkorn University Correspondence: 6531337321@student.chula.ac.th
 Figure 1: An illustration of Partial YaRN.
@@
 Within the Extended Audio Region, the magnitudes of every dimension is scaled to serve as a proxy for attention temperature scaling.
-Despite that the scaling can create a minor discontinuity in the signal, information from other dimensions can help the model disambiguate this.
+Despite that the modality-decoupled scaling can create a minor discontinuity in the signal, information from other dimensions can help the model disambiguate this.
 1 Introduction
@@
 that pair an audio encoder with a text backbone.
-By aligning audio and text in a shared representation, LALMs can leverage the base LLM’s knowledge for complex audio understanding.
+By aligning audio and text into a shared representation space, LALMs can leverage the base LLM’s knowledge for complex audio understanding.
 However, practical use is constrained by short audio context windows: models such as SALMONN (Tang et al., 2024) and Qwen2-Audio (Chu et al., 2024) are typically trained on audio segments of 30s or less, and thus generalize poorly to longer inputs (Guo et al., 2025).
@@
 This potential drawback motivates a more targeted strategy.
-We introduce “Partial YaRN”, an audio-only context extension method that modifies only the audio tokens’ positional encodings.
+We introduce Partial YaRN, a modality-decoupled, audio-only context extension method that modifies only the audio tokens’ positional encodings.
 This design enables a direct comparative study against the whole-context approaches, allowing us to investigate the tradeoff between preserving position encodings of the text modality and maintaining a globally uniform positional space.
-Our experiments on both training-free and fine-tuned audio context extensions reveal that extension-based methods outperforms original models.
-However, the best choice between whole-context and audio-only extensions is often model-dependent without a universally superior solution.
-This suggests that to build truly robust long-audio-context models, we must address the core problem of generalization during the training process.
-Therefore, we extend Partial YaRN into a novel fine-tuning technique called “Virtual Longform Audio Training” (VLAT).
+Our experiments on both training-free and fine-tuned audio context extensions reveal that extension-based methods outperform original models.
+However, the best choice between whole-context and modality-decoupled extensions is often model-dependent without an universally superior solution.
+This suggests that to build truly robust long-audio-capable models, we must address the core problem of generalization during the training process itself.
+Therefore, we extend Partial YaRN into a novel fine-tuning technique called Virtual Longform Audio Training (VLAT).
 Acting as a positional augmentation technique, it simulates a wide range of audio lengths during fine-tuning, teaching the model to generalize beyond the lengths present in the training dataset.
-Through VLAT, we obtain excellent results on longform audio of unseen lengths.
+Through VLAT, we obtain substantially better generalization to longform audio of unseen lengths.
 Our main contributions are as follows:
   • We propose Partial YaRN, a training-free audio-context extension for LALMs that preserves the base language model’s text capabilities by leaving text positions unaltered.
-  • We propose Virtual Longform Audio Training (VLAT), a training strategy that applies Partial YaRN as positional augmentation to simulate diverse audio lengths during training, improving long-context generalization.
-  • We present a comparative study between existing context extension methods designed for unimodal LLMs and our proposed methods for LALMs, analyzing their performance trade-off in both training-free and fine-tuning settings.
+  • We propose Virtual Longform Audio Training (VLAT), a training strategy that extends Partial YaRN into a positional augmentation to simulate diverse audio lengths during training, for improved long-context generalization.
+  • We present a comparative study between existing context extension methods designed for unimodal LLMs and our proposed methods for LALMs, analyzing their performance tradeoff in both training-free and fine-tuning settings.
 2 Preliminary
 2.1 Rotary Positional Encoding
-As Transformer (Vaswani et al., 2017) utilizes attention mechanism (Bahdanau et al., 2014) which is position-invariant, explicit positional information has to be provided for the models to differentiate tokens from different positions.
+As Transformer (Vaswani et al., 2017) utilizes attention mechanism (Bahdanau et al., 2014), which is position-invariant, explicit positional information has to be provided for the models to differentiate tokens from different positions.
 Our study focuses on LALMs that use Rotary Positional Encoding (RoPE) (Su et al., 2024), a widely adopted positional encoding.
@@
 Namely, given a query $\boldsymbol{q_{m}}$ at position $m$, and key $\boldsymbol{k_{n}}$ at position $n$, we have:
-$\boldsymbol{q}_{m}^{\text{T}}\boldsymbol{k}_{n}$ $=(\boldsymbol{R}_{m,\boldsymbol{\theta}}\boldsymbol{q}_{m})^{\text{T}}(\boldsymbol{R}_{n,\boldsymbol{\theta}}\boldsymbol{k}_{n})$ (1)
-$=\boldsymbol{q}_{m}^{\text{T}}\boldsymbol{R}_{n-m,\boldsymbol{\theta}}\boldsymbol{k}_{n}\,$, (2)
+$\boldsymbol{q}_{m}^{\text{T}}\boldsymbol{k}_{n}$ $=(\boldsymbol{R}_{m,\boldsymbol{\theta}}\boldsymbol{q}_{m})^{\text{T}}(\boldsymbol{R}_{n,\boldsymbol{\theta}}\boldsymbol{k}_{n})$
+$=\boldsymbol{q}_{m}^{\text{T}}\boldsymbol{R}_{n-m,\boldsymbol{\theta}}\boldsymbol{k}_{n}\,$,
 where $\boldsymbol{\theta}$ is the set of angular frequencies (step size) of the rotation.
@@
 For example, if originally each positional step is rotated by $\theta_{i}=10^{\circ}$; to double the context length, the frequency (step size) can simply be halved to $5^{\circ}$.
-This way, the model will never have to attend to positions outside of its trained window; resulting in a more stable approach.
+This way, the model will never have to attend to positions outside of its trained window, resulting in a more stable approach.
 YaRN.
@@
 Additionally, YaRN applies temperature scaling to the logits of attention softmax:
-$\operatorname{Attention}(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\operatorname{softmax}\left(\frac{\boldsymbol{Q}^{T}\boldsymbol{K}}{t\sqrt{d_{k}}}\right)\boldsymbol{V},$
+$\operatorname{AS}(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\operatorname{softmax}\left(\frac{\boldsymbol{Q}^{T}\boldsymbol{K}}{t\sqrt{d_{k}}}\right)\boldsymbol{V},$
 where $t$ is the attention temperature.
@@
 Ding et al.
-(2024) propose LongRoPE to account for non-uniformity across RoPE dimensions through a novel evolutionary search algorithm to determine dimension-wise interpolation factors.
+(2024) proposes LongRoPE to account for non-uniformity across RoPE dimensions through a novel evolutionary search algorithm to determine dimension-wise interpolation factors.
 This method is further improved in LongRoPE2 (Shang et al., 2025).
 2.3 Large Audio-language Models and Their Audio Contexts
-One broadly adopted paradigm of large audio-language models (LALMs) is the unified input space (Tang et al., 2024; Chu et al., 2024; Ding et al., 2025; Goel et al., 2025) which involves transforming audio inputs into the textual input space of a base LLM.
+One broadly adopted paradigm of large audio-language models (LALMs) is the unified input space (Tang et al., 2024; Chu et al., 2024; Ding et al., 2025; Ghosh et al., 2025a) which involves transforming audio inputs into the textual input space of a base LLM.
 By sharing the embedding space, the strong text capability and learned knowledge of the base LLM can be leveraged while simultaneously being augmented with audio understanding ability.
-There also exists other LALM paradigms such as cross attention (Kong et al., 2024; Ghosh et al., 2025) where the audio and text modalities are fused through cross-attention modules.
+There also exists other LALM paradigms such as cross attention (Kong et al., 2024; Ghosh et al., 2025b) where the audio and text modalities are fused through cross-attention modules.
 This work focuses on models under the unified-input-space.
@@
 7s for Pengi (Deshmukh et al., 2023); 30s for Qwen2-Audio (Chu et al., 2024)).
-Concurrent to us, Guo et al.
-(2025) introduces FastLongSpeech, an efficient longform LALM that uses Iterative Fusion to compress long audio representation into compact forms.
-Our work, in contrast, investigates an orthogonal direction of adapting the model’s positional encodings rather than the audio embeddings, and therefore is generally applicable to the existing models without requiring costly retraining.
 3 Partial YaRN
-To extend existing unified input LALMs to longer audio, we hypothesize that LALMs already possess a sufficient general understanding of audio and text, but the bottleneck is their unfamiliarity to audio positions beyond the range seen during audio-text training.
+To extend existing unified input LALMs to longer audio, we hypothesize that “LALMs already possess a sufficient general understanding of audio and text, the main bottleneck is instead their unfamiliarity to audio positions beyond the range seen during audio-text training.
 [^3] This differs from whole-context extension in unimodal LLMs, where the total sequence length exceeds the pre-trained context window, causing an out-of-distribution problem.
@@
 However, this alters positional information for all tokens, including text, risking the degradation of the base LLM, which was pretrained solely in text.
-Motivated by the aforementioned hypothesis, we propose Partial YaRN, an audio-only extension method designed to exclusively stretch the audio window.
-By leaving the text’s positional encodings unaltered, this approach aims to extend [^4] The terms extend and stretch are used interchangeably.
+Motivated by the aforementioned hypothesis, we conceive modality-decoupled extension, and propose Partial YaRN, an audio-only extension method designed to exclusively stretch the audio window.
+By leaving the text’s positional encodings unaltered, this approach aims to extend [^4] We use the terms extend and stretch interchangeably.
 audio context while preserving the base model’s text capability.
 3.1 Methodology
-Inspired by PI (Chen et al., 2023) and YaRN (Peng et al., 2024) for whole-context extension in unimodal LLMs, we adapt the interpolation technique to LALMs by applying it exclusively to the audio region of the base language models.
+Inspired by PI (Chen et al., 2023) and YaRN (Peng et al., 2024) for whole-context extension in unimodal LLMs, we adapt the interpolation technique to LALMs in modality-decoupled manner, applying it exclusively to the audio region of the base language models.
 Concretely, assuming a single audio input, let $L_{\text{audio}}$ be the original audio context length, $L_{\text{audio}}^{\prime}$ be the target length, and $p$ be the position of the first audio token.
@@
 where $\oplus$ denotes concatenation in the temporal axis.
-Note that while the final position index is denoted by $L$, the total token capacity now increases to $L+L^{\prime}_{\text{audio}}-L_{\text{audio}}$ to account for the stretched audio window of a higher token frequency.
+Note that while the final position index is denoted by $L$, the total token capacity now increases to $L+L^{\prime}_{\text{audio}}-L_{\text{audio}}$ to account for the stretched audio window with higher token frequency.
 Furthermore, we only partition the RoPE dimensions into two frequency-based groups (instead of three as in the original YaRN formulation): a low frequency group that undergoes pure interpolation, and a high frequency group that undergoes pure extrapolation.
@@
 2) Tuning Efficiency: This simplification reduces the tuning space of the cutoff index from two coupled variables to a single parameter.
-We provide empirical validation and a more in-depth discussion of the two-group partitioning in Section 7.2 and Appendix B respectively. See Figure 1 for an example depiction of Partial YaRN.
+We provide empirical validation and a more in-depth discussion of the two-group partitioning in Section 7.2 and Appendix A respectively. See Figure 1 for an example depiction of Partial YaRN.
 3.2 Hyperparameters
@@
 YODAS2-MCQA
-We source audio from the English subset of YODAS2 (Li et al., 2023b).
+We source audio from the English subset of YODAS2 (Li et al., 2023).
 We segment audio samples into non-overlapping segments of 1, 2, 5, and 10 minutes.
@@
 3) Whole YaRN (Peng et al., 2024): Frequency-based RoPE interpolation with attention scaling, applied to the whole context window. This serves as the primary whole-context baseline.
-These three baselines require no tuning as Vanilla and Whole PI don’t utilize any hyperparameter, and YaRN has predefined cutoffs and a closed-form formula for obtaining temperature.
+These three baselines require no tuning as Vanilla and Whole PI don’t utilize any hyperparameter, and YaRN has predefined cutoffs and a closed-form formula for temperature.
+[^5] Unlike Whole YaRN for unimodal LLMs, where information per token is relatively consistent, compression of audio tokens varied significantly across different LALMs as mentioned.
+This likely influences the optimal hyperparameter choice considerably, as discussed in Appendix E.2.
 Reference models.
-We also report performance of two proprietary models capable of long audio context: GPT-4o (OpenAI, 2024) and Gemini 2.0 Flash (Team et al., 2023), in order to provide a broader perspective of LALMs and to validate the quality of our dataset [^5] The model versions are gpt-4o-mini-2024-07-18 and gemini-2.0-flash-001 respectively.
-Note that Gemini was also used for generating our YODAS2-MCQA dataset..
+We also report performance of two proprietary models capable of long audio context: GPT-4o (OpenAI, 2024) and Gemini 2.0 Flash (Team et al., 2023), in order to provide a broader perspective of LALMs and to validate the quality of our dataset.
+[^6] The model versions are gpt-4o-mini-2024-07-18 and gemini-2.0-flash-001 respectively.
 Vanilla Inference Partial PI Inference
@@
 5 Main Results
-In this section, we comparatively study the effect of whole-context and audio-only context extensions on both the training-free and fine-tuning settings for audio segments up to 10mins.
+In this section, we comparatively study the effect of whole-context and modality-decoupled context extensions on both the training-free and fine-tuning settings for audio segments up to 10mins.
 5.1 Training-free Audio Context Extension
 First, we extend the audio context length without training.
-We report the results in Table 4.1, showing that: no single method—whole-context or audio-only—is universally superior.
+We report the results in Table 4.1, showing that: no single method—whole-context or modality-decoupled—is universally superior.
 Instead, the performance varies between the different models and the degree of extension.
@@
 Extending from observed audio context.
-As previously mentioned in the beginning of this section that we can observe both model’s innate audio contexts to be around 2mins instead of just 30s, here we experiment with extending from this observed audio context of 2mins to the target lengths.
-For both SALMONN and Qwen2-Audio, performance on the long audio improves substantially when the interpolation is anchored from a 2mins context instead of the original 30s.
+As mentioned in the beginning of this section that we can observe both model’s innate audio contexts to be around 2mins instead of just 30s, here we experiment with extending from this observed audio context of 2mins, to the target lengths.
+For both SALMONN and Qwen2-Audio, performance on the long audio improves substantially when the interpolation is anchored from the 2mins context instead of the original 30s.
 Notably on Qwen2-Audio, this strategy boosts Partial YaRN’s performance at 10mins from 28.53% to a 48.00%, an absolute improvement of almost 20%, yielding a final performance 26% higher than the Vanilla baseline.
-This finding suggests that determining the extension by observing the vanilla performance may be a more critical factor than the choice between whole-context or audio-only methods, especially on large extension ratios.
+This finding suggests that determining the extension by observing the vanilla performance may be a more critical factor than the choice between whole-context or modality-decoupled methods, especially on large extension ratios.
 5.2 Fine-tuning for Audio Context Extension
@@
 In the qkvo setting, both Whole YaRN (83.47%) and Partial PI (83.07%) achieve an absolute performance gain of around 19% over the Vanilla baseline (64.93%) at 10mins.
-This highlights the benefit of integrating a context extension method into the fine-tuning process for more better context extension.
-Comparing the whole-context and audio-only approaches, we find that their performance is generally competitive.
-While Whole YaRN triumps when only adapting the query projection weights, suggesting it’s ease of adaptation.
+This highlights the benefit of integrating a context extension method into the fine-tuning process for better context extension.
+Comparing the whole-context and modality-decoupled approaches, we find that their performance is generally competitive.
+While Whole YaRN triumps when only adapting the query projection weights, suggesting its ease of adaptation.
 Their performance converges under standard LoRA practice which adapts qkvo.
@@
 Partial PI 98.00 93.33 83.07
-Table 3: Finetuned audio context extension performance (Accuracy).
-We finetune Qwen2-Audio on YODAS2-MCQA, under two LoRA settings: adapting only q, and adapting qkvo.
-Each model is finetuned and evaluated using the same extension method.
+Table 3: Fine-tuned audio context extension performance (Accuracy).
+We fine-tune Qwen2-Audio on YODAS2-MCQA, under two LoRA settings: adapting only q, and adapting qkvo.
+Each model is fine-tuned and evaluated using the same extension method.
 6 Virtual Longform Audio Training: Generalizing Beyond the Training Data
-Our previous results show that while the training-free methods are beneficial, their effectiveness is often model and audio-length dependent.
-Based on this, we propose to fine-tune LALMs with Partial YaRN repurposed as a positional augmentation technique we call "Virtual Longform Audio Training" (VLAT).
+Our previous results show that while the training-free methods are beneficial, their effectiveness is often model and length dependent.
+Based on this, we turn to address the core generalization problem during the training process, and propose to fine-tune LALMs with Partial YaRN repurposed as a positional augmentation technique we call “Virtual Longform Audio Training” (VLAT).
 6.1 Methodology
@@
 For each sample, we obtain a “virtual” source length, $L_{\text{virt}}$, as the model’s default audio context length times a factor randomly sampled from the range [1, 5, 10, 15, 20, 25]x.
-We then apply Partial YaRN [^6] Whole YaRN is not used here, as we observe that it diverges quickly when training under the VLAT framework.
+We then apply Partial YaRN [^7] Whole YaRN can’t be used here, as we observe that it diverges quickly when training under the VLAT framework.
 to stretch or compress the $L_{\text{virt}}$-long positional window to $L_{\text{data}}$.
@@
 It is also bidirectional, teaching the model through both compressed and stretched contexts.
-Furthermore, VLAT is a targeted, modality-bound method that modifies only the audio tokens, preserving the base LLM’s original textual space.
+Furthermore, VLAT is a targeted, modality-decoupled method that modifies only the audio tokens, preserving the base LLM’s original textual space.
 6.2 Evaluation
-We fine-tune Qwen2-Audio on a 2mins YODAS2-MCQA training set [^7] To prevent data leakage, the train/test splits were performed at the video level, ensuring no source material overlaps between the training set and any of the test sets.
-using two methods: a standard Vanilla fine-tuning and VLAT.
-We employ 10-epoch qkvo LoRA for the adaptations.
+We fine-tune Qwen2-Audio on the 2mins YODAS2-MCQA training set [^8] To prevent data leakage, the train/test splits were performed at the video level, ensuring no source material overlaps between the training set and any of the test sets.
+using two methods: a standard Vanilla fine-tuning, and VLAT.
+We employ 10-epoch qkvo LoRA.
 We then evaluate the models from both training methods under two different extensions at inference: Vanilla and Partial PI.
-See Appendix E.1 for additional detail on VLAT configuration.
+Results are averaged from 3 random seeds.
+See Appendix E.1 for detail on VLAT configuration.
 6.3 Results
@@
 When Partial PI is applied during evaluation, the VLAT-trained model again outperforms its vanilla-trained counterpart, achieving the highest overall 10mins performance of 81.73%.
-This indicates that VLAT and inference-time extension are compatible strategies, and their combination yields the most robust long-context performance.
+This indicates that VLAT and inference-time extension are compatible strategies, and their combination yields the most robust long-audio-context performance.
+We show in Appendix E.1 that VLAT’s performance can be even more enhanced by using a narrower range for virtual length $L_{\text{virt}}$.
 7 Analysis
@@
 The results are presented in Table 7.1.
-We can see that removing either the frequency grouping or the attention temperature generally leads to a large drop in performance, especially on higher extension ratios.
-Both the frequency partitioning and attention temperature scaling are crucial for robustly extending the audio context length of LALMs.
-Their combined application within the Whole and Partial YaRN frameworks generally yield the most effective and stable performance.
 Partial Whole
@@
 Table 4: Ablation of frequency grouping and attention temperature components (Accuracy). Removing either the frequency grouping or the attention temperature leads to a large drop in performance.
+We can see that removing either the frequency grouping or the attention temperature generally leads to a large drop in performance, especially on higher extension ratios.
+Both the frequency partitioning and attention temperature scaling are crucial for robustly extending the audio context length of LALMs.
+Their combined application within the Whole and Partial YaRN frameworks generally yield the most effective and stable performance.
 7.2 Partial YaRN with Three-group Frequency Partitioning
-For the Partial YaRN method, we modify the original 3-group frequency partition from original YaRN to a 2-group method as discussed in Section 3.1 and Appendix B.
+For the Partial YaRN method, we modify the original 3-group frequency partition from original YaRN to a 2-group method as discussed in Section 3.1 and Appendix A.
 Here we conduct a direct empirical comparison between the 2-group and 3-group approaches.
@@
 This is most evident with the Qwen2-Audio model, where the 3-group method suffers a large performance drop on 5mins and 10mins audio, with an accuracy gap of 28.53% and 12.00% respectively.
-Overall, this result validates the effectiveness and reliability of our proposed 2-group scheme used in Partial YaRN.
-8 Conclusion
+Overall, this result validates the effectiveness and reliability of our proposed 2-group scheme.
+8 Related Work
+This section discusses additional related works that aren’t addressed in Section 2.
+Positional encodings.
+Absolute positional encoding was originally proposed in Vaswani et al.
+(2017).
+It utilizes sinusoidal waves to uniquely represent absolute position of each token.
+Later, (Devlin et al., 2019; Lan et al., 2020; Clark et al., 2020) use trainable vectors for representing absolute positions, allowing the models to learn the positional encoding for themselves.
+However, it cannot extrapolate to sequences longer than the maximum length it was trained on.
+Shaw et al.
+(2018) encodes the relative distances between the tokens instead of absolute positions.
+Dai et al.
+(2019) introduces additional trainable position-agnostic components to distinguish between content-based and location-based attention behaviors.
+Raffel et al.
+(2020) groups the relative distances into buckets, and associate each bucket with a learnable scalar for adding to attention logits.
+TUPE (Ke et al., 2021) uses separate linear projections for the token and positional encodings.
+ALiBi (Press et al., 2022) simply adds linear biases with a learnable slope into the attention scores, imposing a strong locality preference.
+Several later works extend RoPE for multimodal modeling, such as Multimodal RoPE Bai et al.
+(2025), VideoRoPE (Wei et al., 2025), and VRoPE (Liu et al., 2025b).
+Large video-language models.
+Video input is similar to the audio in that they’re both naturally continuous and can be of arbitrary lengths.
+Many Large Video-Language Models (LVLMs) (Zhang et al., 2023b; Maaz et al., 2024; Lin et al., 2024; Jin et al., 2024; Ren et al., 2024), adopt architectural paradigm similar to the unified input space LALMs we studied, where video frames are embedded and projected into a base LLM’s input space.
+However, the context length problem is even more exacerbated in the video domain due to the high dimensionality and large number of tokens needed to represent even short videos.
+To improve efficiency, the LVLM community has largely focused on input compression techniques like query aggregation (Zhang et al., 2023b; Li et al., 2025a; Song et al., 2024), frame pooling (Luo et al., 2023; Maaz et al., 2024), and feature merging (Weng et al., 2024).
+These strategies primarily aim to reduce the number of tokens before they enter the base LLM.
+Our work offers a complementary perspective: instead of altering the input, we manipulate the positional space within the model to accommodate a longer sequence, similar to Zhang et al.
+(2025) in LVLMs.
+The principles of Partial YaRN and VLAT are likely applicable to the video domain as well, hopefully providing a distinct path to achieve long-video-context understanding and generative control in synergy with existing methods.
+Long-audio LALMs.
+Concurrent works have proposed LALMs that have inherently long audio-context.
+Audio Flamingo 3 (Ghosh et al., 2025a) is capable of processing audio up to 10mins due long-audio training.
+Voxtral (Liu et al., 2025a) can handle up to 40mins due to the long 32k context of its LLM backbone.
+FastLongSpeech (Guo et al., 2025) proposes Iterative Fusion which incrementally merges the audio representation based on their information density.
+This compressed representation facilitates efficient processing of long audio.
+Qwen3-Omni (Xu et al., 2025) is a multimodal model that can accept text, video, and audio as input.
+It employs a custom Audio Transformer which is trained in an encoder-decoder fashion similar to the Whisper model.
+Qwen3-Omni is also capable of processing audio of up to 40mins.
+While these models achieve impressive long-audio performance, they do so by designing new architectures, curating large volumes of long-audio data, and undertaking extensive training.
+Our work investigates an orthogonal yet complementary direction of efficiently extending the context of existing LALMs.
+Rather than requiring costly data acquisition and retraining, we provide a lightweight, architecture-agnostic "drop-in" enhancement (Partial YaRN) and an efficient fine-tuning strategy (VLAT), offering a practical and accessible approaches for improving the long-audio capabilities of a wide range of existing models.
+9 Conclusion and Future Work
 This work addressed the challenge of extending audio context windows of LALMs by studying the application of LLM’s whole-context extension methods, and proposing a training-free audio-only extension Partial YaRN and the VLAT training strategy.
-We demonstrate the effectiveness of extension-based methods over non-extension baselines across various long-audio settings.
+We demonstrated the effectiveness of extension-based methods over non-extension baselines across various long-audio settings.
 Leveraging our finding that SALMONN and Qwen2-Audio retain consistent performance up to 2mins audio window, we achieved an even stronger performance by extending from the 2mins window instead of 30s.
-Later, we showed that fine-tuning through VLAT helps the models to generalize to audio contexts far exceeding the length of the training data, offering a robust and data-efficient pathway to develop LALMs with better longform understanding.
-Future work could explore a similar modality-bound extension strategy for video models.
+Later, we showed that fine-tuning through VLAT allows the models to generalize to audio contexts far exceeding the lengths of the training data, offering a robust and data-efficient pathway to develop LALMs with better longform understanding.
+Future work could apply Partial YaRN and VLAT to vision models.
+Beyond content understanding, these approaches could also be explored for controllable generation in multimodal models, such as multi-scale image synthesis, or even more finegrained decoupling for non-uniform speed control throughout audio or video generation.
 Limitations
 Evaluation scope.
-Our experiments are conducted exclusively on the multiple-choice question answering (MCQA) tasks.
+Our experiments are conducted exclusively on the multiple-choice question answering tasks.
 We made this choice to ensure a precise and unambiguous measurement of the model performance under different audio lengths.
@@
 Consequently, while our results validate the model’s ability to retrieve information from audio, they do not fully test the importance of text preservation of the audio-only methods like Partial YaRN.
-Furthermore, our study does not assess performance across other long-audio task formats found in benchmarks such as (Guo et al., 2025; Ahia et al., 2025; Goel et al., 2025).
+Furthermore, our study does not assess the performance across other long-audio task formats found in benchmarks such as Guo et al.
+(2025), Ahia et al.
+(2025), or Ghosh et al.
+(2025a).
+Reliance on synthetic data.
+Our study utilizes YODAS2-MCQA, a custom dataset synthesized from the YODAS2 corpus using Gemini 2.0 Flash.
+While we carefully steered the synthesis process and validated a subset of the data, it’s still possible for the synthesized samples to carry noise, repetition, and hallucinations that may affect the fidelity of our performance estimates compared to real-world scenarios.
 Hyperparameter tuning burden.
-Unlike the original YaRN which has generally good predefined cutoff dimensions and a closed-form formula for determining the attention temperature, our proposed Partial YaRN requires tuning of these two hyperparameters, thereby introducing additional computational burden.
+Unlike the original YaRN which has generally good predefined cutoff dimensions and a closed-form formula for determining the attention temperature, our proposed Partial YaRN requires tuning of two hyperparameters, thereby introducing additional computational burden.
 Acknowledgments
-YC would like to thank Ekapol Chuangsuwanich for insightful discussion on the earlier version of this work. He also thanks Thanapat Trachu for valuable guidance on the illustration of Partial YaRN.
-Appendix A Extended Related Works
-This section discusses additional related works that aren’t addressed in Section 2.
-Positional encodings.
-Conceived along with the introduction of transformers (Vaswani et al., 2017), the original absolute positional encoding utilizes sinusoidal waves of varying frequencies to uniquely represent absolute position of the tokens, and is incorporated into the model by adding directly into the input sequence.
-Later works (Devlin et al., 2019; Lan et al., 2020; Clark et al., 2020) propose to use trainable vectors for representing absolute positional information, where token at each position is summed with a learned vector, allowing the models to learn the optimal positional encoding for themselves.
-However, it cannot extrapolate to sequences longer than the maximum length it was trained on.
-Shaw et al.
-(2018) introduces to encode the relative position of the token instead, where the relative distance between the query and key is captured and injected into every attention layer.
-Dai et al.
-(2019) modifies the decomposed form of the attention equation, and introduces additional trainable location-agnostic key vectors to distinguish between content-based and location-based querying behaviors.
-Raffel et al.
-(2020) groups the relative distances into buckets, and associate each bucket with a learnable scalar for adding into the attention logits.
-Simplifying the encoding scheme and reducing trainable parameters.
-TUPE (Ke et al., 2021) uses separate linear projections for the word and positional information.
-ALiBi (Press et al., 2022) simply adds linear biases with a learnable slope into the attention scores, the bias grows linearly with the relative distance between the tokens.
-This method generally imposes a strong locality preference by encouraging models to attend more to the nearby tokens, while discouraging them from attending to tokens that are farther away.
-Several later works extend RoPE for multimodal modeling (e.g.
-Multimodal RoPE Bai et al.
-(2025), VideoRoPE (Wei et al., 2025), VRoPE (Liu et al., 2025)).
-Large video-language models.
-Video input is similar to the audio modality in that they’re both naturally continuous and can be of arbitrary length.
-Many Large Video-Language Models (LVLMs), such as Video-LLaMA (Zhang et al., 2023b) and Video-ChatGPT (Maaz et al., 2024), adopt architectural paradigm similar to the unified input space LALMs we study, where video frames are embedded and projected into a base LLM’s input space.
-However, the context length problem is even more exacerbated in the video domain due to the high dimensionality and large number of tokens needed to represent even short videos.
-To improve efficiency, the LVLM community has largely focused on input compression techniques like query aggregation (Zhang et al., 2023b; Li et al., 2023a; Song et al., 2024), frame pooling (Luo et al., 2023; Maaz et al., 2024), and feature merging (Weng et al., 2024).
-These strategies primarily aim to reduce the number of tokens before they enter the LLM.
-Our work offers a complementary perspective: instead of altering the input, we manipulate the positional space within the model to accommodate a longer sequence.
-The principles of Partial YaRN and Virtual Longform Audio Training could therefore be potentially applicable to the video domain as well, hopefully providing an orthogonal path to achieving long-context understanding in synergy with existing compression methods.
-Appendix B Elaboration on Two-group Partitioning of Partial YaRN
+YC would like to thank Ekapol Chuangsuwanich for insightful discussion on the earlier version of this work.
+He also thanks Thanapat Trachu for valuable guidance on the illustration of Partial YaRN, and Konpat Preechakul for illuminating discussion on extensions of this work to the vision domain.
+We gratefully acknowledge the compute resources provided by Hatari NEXT, on which the experiments for this research were run.
+Appendix A Elaboration on Two-group Partitioning of Partial YaRN
 The decision to simplify the original three-group frequency partition of YaRN (Peng et al., 2024) into a two-group system for Partial YaRN is a deliberate design choice aimed at ensuring positional consistency across the entire extended audio stream.
@@
 This ensures that the model maintains a consistent positional understanding across all dimensions for the entirety of the audio input, avoiding the representational bias that partial coverage would introduce.
-See Appendix 7.2 for direct performance comparison between 2 and 3-group frequency partitioning versions of Partial YaRN.
-Appendix C Implementation of Partial YaRN via position_ids Manipulation
+See Section 7.2 for direct performance comparison between 2 and 3-group frequency partitioning versions of Partial YaRN.
+Appendix B Implementation of Partial YaRN via position_ids Manipulation
 A key implementation detail of our work is how we apply the positional stretching to only a subsection of the input sequence.
-While the original YaRN implementation modifies the RoPE frequency variables ($\boldsymbol{\theta}$), this approach isn’t straightforward for a partial modification.
+While the original YaRN implementation modifies RoPE’s frequencies ($\boldsymbol{\theta}$), this approach isn’t straightforward for a partial modification.
 Instead, our implementation of Partial YaRN directly manipulates the position_ids and multiply it with the unmodified frequencies.
@@
 Therefore, the cost of enabling partial context extension is a small, fixed precomputation that does not impact the per-token generation latency.
+Appendix C On Deriving Closed-Form Hyperparameter Formulae
+A practical extension of this work is to derive closed-form formulae for determining the optimal cutoff dimension index and attention temperature for Partial YaRN akin to the solution provided in YaRN for unimodal LLMs (Peng et al., 2024), which would eliminate the cost of hyperparameter searches.
+We leave this for future work due to a key challenge not present in unimodal LLMs: the wide variance in audio representation lengths across different LALMs.
+Unlike the relatively consistent token-to-information ratio in text across modern tokenizers, LALMs compress audio into discrete tokens at vastly different rates.
+For instance, a 30-second audio clip gets compressed into 88 tokens in SALMONN but 750 tokens in Qwen2-Audio.
+As our sensitivity analysis in Appendix E.2 shows, this disparity strongly influences the optimal hyperparameter choice, making closed-form formula non-trivial to derive.
 Appendix D Additional Experimental Details
@@
 We use LoRA (Hu et al., 2022) of rank 8 for fine-tuning through HuggingFace’s PEFT library (Mangrulkar et al., 2022).
-With alpha set to 16, and a small LoRA dropout of 0.5%.
+With alpha set to 16, and a small LoRA dropout of 5%.
 Models are trained with AdamW optimizer (Loshchilov and Hutter, 2019), learning rate is set to 5e-5, a batch size of 8 samples, and gradient norm clipped to 1.0 (Pascanu et al., 2013).
@@
 See Appendix E.2 for tuning results and sensitivity.
-D.3 Dataset Construction
-We specifically use a portion the en129888https://huggingface.co/datasets/espnet/yodas2/tree/main/data/en129 subset of the YODAS2 dataset (Li et al., 2023b) (CC-BY-3.0) through HuggingFace’s Datasets library (Lhoest et al., 2021).
-The dataset is then split into train, validation, and test sets.
-We guarantee no test leakage in a way that if a dialogue is in the test set, then all the other dialogues from the same video are enforced to be in the test set as well.
-YODAS2 MCQA inference prompt Your answer MUST be in the format of "(NUMBER) STATEMENT".
+Figure 2: Distributions of audio segment lengths in the YODAS2-MCQA dataset, shown with logarithmic counts for improved visibility of the long-tail portions.
+Figure 3: Distributions of correct answer choices for all question-answer pairs in the YODAS2-MCQA dataset.
+YODAS2-MCQA inference prompt Your answer MUST be in the format of "(NUMBER) STATEMENT".
 For example, if the answer was (4) A pen, you would ONLY output "(4) A pen".
@@
 Figure 4: Prompt used for inferencing on YODAS2-MCQA.
+D.3 Dataset Construction
+Our dataset is derived from a portion of the en129999https://huggingface.co/datasets/espnet/yodas2/tree/main/data/en129 subset of YODAS2 (Li et al., 2023), a large-scale, multilingual corpus of YouTube-derived audio and speech data.
+The YODAS2 dataset is made available under the CC-BY-3.0 license.
+We accessed and processed the data using HuggingFace’s Datasets library (Lhoest et al., 2021).
+To prevent video source leakage in the VLAT experiment, train-test split was performed at the video level; all dialogue segments from a single video were exclusively assigned to either the training set (2mins), or the test sets.
 To obtain specific audio length, we chunk the long audio in the dataset to 1, 2, 5, and 10 minutes.
-Last, incomplete chunk of each long-audio is also included as well.
-Each of the chunk will be used to generate 5 MCQA pairs via Gemini 2.0 Flash (Team et al., 2023).
-We provide the full MCQA generation prompt in Figure 7.
+Any remaining, incomplete chunk of each long-audio is also retained.
+Subsequently, we employed Gemini 2.0 Flash (Team et al., 2023) to generate five unique multiple-choice question-answer (MCQA) pairs for each audio chunk.
+The complete prompt utilized for this generation process is provided in Figure 9.
+The resulting dataset, which we have named YODAS2-MCQA, is publicly available at: https://huggingface.co/datasets/yophis/yodas2-mcqa.
+For a more granular analysis of the YODAS2-MCQA dataset, we provide visualizations of the distributions of key characteristics.
+The distributions of audio segment lengths are presented in Figure 2.
+Additionally, the distribution of correct answers across the multiple-choice options is presented in Figure 3.
 Appendix E Further Experimental Results
@@
 We report the result in Table 6.
-YODAS2-MCQA
-Virtual Length Strategy 2 mins 5 mins 10 mins
-Default 96.87 91.91 81.73
-100 Sampling Points 95.91 89.25 79.78
-1000 Sampling Points 96.18 87.69 79.60
-Limited Range 98.00 93.55 89.20
-Fixed to 20.0x 89.60 68.71 50.31
-Table 6: VLAT performance with different sampling strategies. All models are trained on 2mins audio and evaluated with Partial PI inference. The results are averaged across three random seeds.
 Analysis.
@@
 Given this uncertainty, we retained the Default strategy for our main experiments in Section 6.
-Its factor range intuitively and completely covers the lengths seen during evaluation, making it a more principled and safer choice.
+Its factors ranges intuitively and completely covers the lengths seen during evaluation, making it a more principled and safer choice.
+YODAS2-MCQA
+Virtual Length Strategy 2 mins 5 mins 10 mins
+Default 96.87 91.91 81.73
+100 sampling points 95.91 89.25 79.78
+1000 sampling points 96.18 87.69 79.60
+Limited range 98.00 93.55 89.20
+Fixed to 20.0x 89.60 68.71 50.31
+Table 6: VLAT performance with different sampling strategies. All models are trained on 2mins audio and evaluated with Partial PI inference. The results are averaged across three random seeds.
 E.2 Hyperparameter Sensitivity of Partial YaRN
@@
 Figure 8: Hyperparameter Sensitivity of Partial YaRN on SALMONN (extending the observed 2mins audio context).
-YODAS2 MCQA generation prompt You are an expert in creating high-quality Multiple-Choice Question Answering (MCQA) datasets for speech understanding.
+YODAS2-MCQA generation prompt You are an expert in creating high-quality Multiple-Choice Question Answering (MCQA) datasets for speech understanding.
 Your task is to generate QA pairs based on audio segments.